What is the minimum of the function x^2+4x+2?
Answer by stanbon(75887)

minimum of the function x^2+4x+2?
----------------
There are many ways to do this problem, depending
on where you are in your studies. One way is to
find x=-b/2a
For your problem a=1,b=4, so x=-4/2=-2
Then find f(-2)=(-2)^2+4(-2)+2= -2
The minimum is at (-2,-2)
Cheers,
Stan H.